a) Suppose that 𝑋1,𝑋2,...,𝑋𝑛 is a random sample from a distribution with

probability density function:

𝑓(π‘₯;πœ†)={

1

6πœ†4π‘₯3π‘’βˆ’π‘₯

πœ†, 𝑖𝑓 π‘₯ β‰₯0 π‘Žπ‘›π‘‘ πœ†>0

0, 𝑒𝑙𝑠𝑒 π‘€β„Žπ‘’π‘Ÿπ‘’

i) Show whether or not 𝑓(π‘₯;πœ†) belongs to the 1-parameter exponential family.

(5)

ii) Derive the Maximum likelihood estimator of πœ†. (8)

iii) Suppose that 𝑛=200, βˆ‘ π‘₯𝑖200𝑖=1 =20,βˆ‘ π‘₯𝑖2200𝑖=1 =100, and βˆ‘ π‘₯𝑖3200𝑖=1 =250.

Determine the Maximum likelihood estimate of πœ†. (2)


1
Expert's answer
2021-10-19T11:29:00-0400

i)

"\int_0^\infty x^3e^{-x\lambda}dx=-\frac{1}{\lambda}x^3e^{-x\lambda}|_0^\infty+\frac{3}{\lambda}\int_0^\infty x^2e^{-x\lambda}dx=0-\frac{6}{\lambda^2}x^2e^{-x\lambda}|_0^\infty+\frac{6}{\lambda^2}\int_0^\infty xe^{-x\lambda}dx=0-\frac{6}{\lambda ^3}xe^{-x\lambda}|_0^\infty+\frac{6}{\lambda^3}\int_0^\infty e^{-x\lambda}dx=-\frac{6}{\lambda ^4}e^{-x\lambda}|_0^\infty=0+\frac{6}{\lambda ^4}=\frac{6}{\lambda ^4}"

Therefore "\int_0^\infty(\frac{1}{6}\lambda^4 x^3 e^{-x\lambda})dx=1"

I.e. "p(x.\lambda)=\frac{1}{6}\lambda^4 x^3 e^{-x\lambda},x>0" is exponential 1 parametric distribution.

2) Maximum likehood estimator of "\lambda" :

"L(\lambda)=p(x_1,...,x_n|\lambda)=\prod_{i=1}^np(x_i\lambda)=\frac{\lambda^{4n}}{6^n}(\prod_{i=1}^{n}x_i)^3e^{-\lambda\sum_1^nx_i}"

"lnL(\lambda)=4n\cdot ln(\lambda)-n\cdot ln(6)+3ln" "(\prod_{i=1}^{n}x_i)"- "\lambda\sum_1^nx_i" ;

"(lnL(\lambda))'=\frac{4n}{\lambda}-" "\sum_{i=1}^nx_i=0;"

"\lambda=\frac{4n}{\sum_{i=1}^nx_i}=\frac{4}{\overline x}" - the likehood estimator.

3)

"\hat\lambda=\frac{4\cdot 200}{\sum_{i=1}^{200}x_i}=\frac{800}{20}=40"